# Shortcut Learning in In-Context Learning: A Survey

### v1

@@ Abstract
Shortcut learning refers to the phenomenon where models employ simple, non-robust decision rules in practical tasks, which hinders their generalization and robustness. With the rapid development of large language models (LLMs) in recent years, an increasing number of studies have shown the impact of shortcut learning on LLMs. This paper provides a novel perspective to review relevant research on shortcut learning in In-Context Learning (ICL). It conducts a detailed exploration of the types of shortcuts in ICL tasks, their causes, available benchmarks, and strategies for mitigating shortcuts. Based on corresponding observations, it summarizes the unresolved issues in existing research and attempts to outline the future research landscape of shortcut learning.
Shortcut Learning in In-Context Learning: A Survey
Rui Song1,2, Yingji Li1,2, Fausto Giunchiglia1,3 Hao Xu1, 1College of Computer Science and Technology, Jilin University, Changchun, China, 2Key Laboratory of Symbol Computation and Knowledge Engineering of Ministry of Education, Jilin University, China, 3Department of Information Engineering and Computer Science, University of Trento, Italy Correspondence: xuhao@jlu.edu.cn
1 Introduction
@@
In recent years, Large Language Models (LLMs) have emerged as a hotly pursued research direction, with the advent of major language models such as T5 Raffel et al. (2020), LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2023), GPT-3 Brown et al. (2020), Qwen2 Yang et al. (2024a), and GLM Zeng et al. (2023). LLMs demonstrate the ability of In-Context Learning (ICL), meaning they can learn from several demonstration examples within a given context Dong et al. (2022). However, an increasing number of studies have shown that LLMs are susceptible to shortcut learning in ICL and perform poorly in many aspects Du et al. (2024).
@@
Shortcuts refer to decision rules that perform well on standard benchmarks but fail to transfer to more challenging testing conditions. Shortcut learning is the process by which a model exploits these decision rules during execution, leading to robust output results Geirhos et al. (2020). In the literature, shortcut learning is also referred to as superficial/spurious correlations Zhou et al. (2024c), the Clever Hans effect Lapuschkin et al. (2019), and various types of biases 111In our survey, bias refers to systematic errors within LLMs, rather than cultural or social biases that are studied in the field of LLM safety and fairness Zheng et al. (2024)., such as learning bias Du et al. (2024), label bias Fei et al. (2023), selection biases Wei et al. (2024), co-occurrence bias Kang and Choi (2023), and so forth. In practice, if LLMs establish shortcut ‘FlowersPosition’ based on given samples in Figure 1, predictions may fail when confronted with ‘Negative’ samples. Apart from this mapping from lexicon to specific labels, shortcuts also exist widely in various forms across different NLU tasks. For example, in Natural Language Inference (NLI) Zhou and Bansal (2020), Question-Answering (QA) Sen and Saffari (2020), and Reading Comprehension (RC) Lai et al. (2021), there is a preference for overlapping parts of two input branches. These unhealthy shortcuts have been widely proven to be harmful to the robustness, generalization, and fairness, and may lead to hallucinations in LLMs Li et al. (2023b); Ji et al. (2023).
@@
Given the rapid development of LLMs and the impact of shortcut learning on them, our survey aims to systematically summarize, discuss, and consolidate the existing research, and to sensitize the community to the current advancements. Different from previous reviews on shortcut learning Ho et al. (2022); Du et al. (2024); Dogra et al. (2024); Bihani and Rayz (2024), our survey focuses on LLMs, while prior research primarily centered on Lightweight Pre-trained Language Models (LPLMs) such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019). Considering the differences in the training and application of LLMs, we deem it necessary to reassess the types of shortcuts, the causes of shortcut learning, benchmarks, and mitigation strategies, and to highlight the unresolved issues surrounding shortcut learning in LLMs. The organization of sections is as follows.
@@ 2 Types of Shortcuts
Compared to Du et al. (2024), we extend and summarize various types of shortcuts based on the characteristics of NLU tasks in LLMs. Specifically, we classify shortcuts into instinctive shortcuts and acquired shortcuts, based on whether LLMs rely on prompts to establish superficial correlations with labels or answers. But in reality, we believe that instinctive shortcuts can influence acquired shortcuts to some degree. In other words, instinctive shortcuts may facilitate LLMs in acquiring shortcuts from demonstrations Pan et al. (2023).
@@ 2.1 Instinctive Shortcuts
Instinctive shortcuts refer to the inherent preferences of LLMs for texts or patterns outside the examples provided in a given context. These preferences are not influenced by the contextual demonstrations but hinder the effective learning of the input-label relationships presented in those demonstrations. To avoid naming ambiguity, we have adopted the label bias definition proposed by Fei et al. (2023), which has been widely discussed.
@@
Vanilla-label bias describes the model’s uncontextual preference for predicting certain label names. Zhao et al. (2021); Zheng et al. (2024) point out that regardless of the prompt, LLMs tend to assign higher probabilities to specific labels. Holtzman et al. (2021) demonstrates the existence of Surface Form Competition among different words, leading to a significant decline in prediction performance when low-frequency words are the correct answers. Similarly McCoy et al. (2023); Kang and Choi (2023) also confirm the preference of LLMs for high-frequency labels. Pan et al. (2023) show that replacing labels with meaningless words, such as ‘@#$’, leads to a decline in performance compared to using meaningful labels like ‘ABC’.
@@
Context-label bias reveals the influence of context prompt, including factors such as sequence and format, on prediction outcomes. For example, subtle adjustments to the prompt format can lead to high instability in model performance, such as changing ‘Passage:<text>’ to ‘Passage <text>’ Sclar et al. (2024). Additionally, LLMs are also highly sensitive to term frequencies and task probabilities Razeghi et al. (2022). For instance, when asked to decode shift ciphers with various shift levels, GPT-4 scores 50% or above for the three most common shift numbers (1, 3, and 13), but scores below 3% for all other shift levels represented by different numbers McCoy et al. (2023). Moreover, LLMs are also highly sensitive to the order of inputs, whether it be the order of demonstrations or the order of candidate answer options. For example, LLMs are thought to favor options presented at a particular ranking position, such as first or last Wang et al. (2024); Pezeshkpour and Hruschka (2024); Zheng et al. (2024); Wei et al. (2024). While Lu et al. (2022) demonstrates that the sequence of presented samples can lead to a vast difference in performance, ranging from nearly achieving the current state-of-the-art to merely performing at the level of random guessing.
@@
Domain-label bias describes the dependence on prior knowledge related to the task. Research has indicate that the in-domain words randomly sampled from the dataset are heavily biased towards the domain labels in ICL Fei et al. (2023). Accordingly, label words also often imply specific task types Si et al. (2023a). For example, the word ‘positive’ frequently appears in connection with sentiment classification tasks. If ‘positive’ were replaced with a meaningless symbol like ‘1’, LLMs would find it difficult to discern whether the task is sentiment classification or topic classification. This leads to LLMs relying on their pretrained demonstration semantic priors in ICL predictions, rather than learning from the input-label relationships presented in demonstrations Jang et al. (2024). This prior can also be observed in QA tasks, where LLMs are able to make correct predictions even without being given options Balepur et al. (2024).
@@
Apart from the semantic priors between texts and labels, entity knowledge within LLMs can also lead to bias Wang et al. (2022). For instance, LLMs may assume that ‘Bill Gates’ is the founder of ‘Microsoft’, even when given the context ‘Bill Gates went to Microsoft Building 99’ where the relationship between ‘Bill Gates’ and ‘Microsoft’ is visitor. Similarly, in relation extraction tasks where the intended relationship to be extracted is unrelated to the existing relationships in the input, LLMs may still extract these unrelated relationships Zhang et al. (2024b). In multi-hop QA tasks, LLMs tend to directly provide answers by leveraging potential associations between the head entity and the entity to be predicted, rather than engaging in correct reasoning paths Li et al. (2024a); Ju et al. (2024).
@@
Reason bias is a newly summarized shortcut type in this paper compared with Fei et al. (2023). It indicates that LLMs tend to get answers directly from the provided examples rather than reasoning the underlying patterns Ali et al. (2024), especially in complex reasoning and summarization tasks involving long contexts Li et al. (2024b); Wu et al. (2024); Yang et al. (2024c). At first glance, reason bias may seem to rely on demonstrations, but essentially, it represents a tendency of LLMs to take shortcuts when solving complex problems. Therefore, we categorize it as an instinctive shortcut.
@@ 2.2 Acquired Shortcuts
Lexicon. When spurious correlation are established between certain lexical features and specific labels, the corresponding lexicon may then become a shortcut. Arbitrary text can serve as lexicon shortcuts, including letters, signs, numbers, negation words, adverbs of degree, or even recurring sentences Tang et al. (2023). Among them, negation words have garnered the most attention from researchers, as LLMs often provide contradictory predictions due to these words Sun et al. (2024); Yuan et al. (2024); Zhou et al. (2024d) But studies have indicated that LLMs are more likely to rely on n-grams and content words rather than others, such as stop words Si et al. (2023b); Pacchiardi et al. (2024).
@@
Concept. Concept shortcut refers to the phenomenon where the concepts of certain texts in the demonstrations are disproportionately associated with the labels Zhou et al. (2024c). For example, if sentences containing cities are associated with negative emotions and samples containing countries are also associated with negative emotions, when LLMs encounter new city words in the test sample, they may predict it as having negative emotions Zhou et al. (2024d).
@@
Overlap. Overlap shortcuts often emerge in tasks that involve two branches, such as NLI and QA. In NLI, overlap between the premise and the hypothesis can be a strong indicator of Entailment Sun et al. (2024); Yuan et al. (2024). In extractive QA, words that appear in both the question and the background may be used as anchors. LLMs tend to extract answers around these anchors to reduce the error rate Levy et al. (2023).
@@
Position. LLMs use positional information as a basis for prediction. For example, if the answer is located at the beginning of the text, the model is less likely to make an error Levy et al. (2023). LLMs can also briefly read the opening segment to predict the answer without needing to read the entire document Li et al. (2024b). Conversely, Yuan et al. (2024) show that adding content unrelated to the answer at the beginning of a sentence can degrade the prediction performance of LLMs. Additionally, Zhao et al. (2021) also observe that answers closer to the end of the prompt are more likely to be repeated by the model.
@@
Text Style. The writing style of the text is also a possible shortcut Qi et al. (2021). Tang et al. (2023); Yuan et al. (2024) demonstrate that shifts in presentation style can have an impact on performance, such as when standard English is rewritten into Shakespearean English. LLMs can also capture the correlation between style and labels, leading to a decline in the generalization ability of ICL Zhou et al. (2024d).
@@
Group Dynamics. It refers to the interactions, relationships, and influences among members within a group Cartwright and Zander (1968). Similarly, in the decision-making process, LLMs may evaluate a choice based on other available options. For instance, if other options are even numbers, the model may tend to lean towards an odd number choice Balepur et al. (2024). This type of shortcut can also be associated with the A-Not-B error in the field of cognitive science Popick et al. (2011). If ‘A’ appears more frequently in the demonstrations, the model is more inclined to predict ‘A’, even if ‘B’ is correct Gupta et al. (2023); Han et al. (2024).
@@ 3.1 LLMs training
Instinctive shortcuts reveal the inherent biases of LLMs. The causes of inherent biases are diverse and closely related to the training process of LLMs, involving factors such as data distribution and instruction tuning. This makes it difficult for LLMs to fully overcome the prediction biases acquired from pre-training data and to consider all contextual information equally during the ICL process Kossen et al. (2024).
Pretraining. Some studies suggest that instinctive shortcuts originate from the data distribution during the pretraining process Razeghi et al. (2022); McCoy et al. (2023), which leads to a bias in the predictions of LLMs towards high-frequency and frequently co-occurred words Kang and Choi (2023); Ju et al. (2024). Rare behavioral patterns in the training data, such as converse relations, can also confuse LLMs, causing them to rely on shortcuts when making judgments Qi et al. (2023). Furthermore, Zhou et al. (2024d) utilize SHAP analysis Lundberg and Lee (2017) to confirm that high-frequency words that have spurious correlations with a certain label contribute more significantly to the model’s predictions for that label. Additionally, if the frequency of a particular LLMs task is low, its prediction results will also be inferior compared to those of high-frequency tasks McCoy et al. (2023). Based on the observations mentioned above, Han et al. (2022); Fei et al. (2023) implicitly assume that there is a shift in the label marginal of the ICL predictive distribution, and Jiang et al. (2023) theoretically proves that the emergence of shortcuts stems from label shifting caused by the contextual label marginal typically deviating from the true label distribution.
Instruction Tuning. Another perspective holds that the instinctive shortcuts in LLMs originate from the process of instruction tuning, during which the LLMs learn spurious correlations within the instructions Sun et al. (2024). In complex multitasking scenarios, LLMs can also learn more intricate associations between tasks, features, and labels during the fine-tuning process. As a result, LLMs can exploit these spurious correlations simply by being provided with useful task information Pan et al. (2023).

### v2

@@ Abstract
Shortcut learning refers to the phenomenon where models employ simple, non-robust decision rules in practical tasks, which hinders their generalization and robustness. With the rapid development of Large Language Models (LLMs) in recent years, an increasing number of studies have shown the impact of shortcut learning on LLMs. This paper provides a novel perspective to review relevant research on shortcut learning in In-Context Learning (ICL). It conducts a detailed exploration of the types of shortcuts in ICL tasks, their causes, available benchmarks, and strategies for mitigating shortcuts. Based on the corresponding observations, we summarize the unresolved issues in existing research and attempts to outline the future research landscape of shortcut learning.
keywords:
Shortcut Learning, Large Language Models, In-Context Learning, Natural Language Processing, Literature Review1 Introduction
@@
In recent years, Large Language Models (LLMs) have emerged as a hotly pursued research direction, with the advent of major language models such as T5 [1], LLaMA [2], PaLM [3], GPT-3 [4], Qwen2 [5], and GLM [6]. LLMs demonstrate the ability of In-Context Learning (ICL), meaning they can learn from several demonstration examples within a given context without fine-tuning [7]. ICL has been widely confirmed as an emerging capability of LLMs and has led to new usage paradigms for LLMs [8, 9]. However, an increasing number of studies have shown that LLMs are susceptible to shortcut learning in ICL and perform poorly in many aspects [10].
@@
Shortcuts refer to decision rules that perform well on standard benchmarks but fail to transfer to more challenging testing conditions. Shortcut learning is the process by which a model exploits these decision rules during execution, leading to non-robust output results [11]. In the literature, shortcut learning is also referred to as superficial/spurious correlations [12], the Clever Hans effect [13], and various types of biases111In our survey, bias refers to systematic errors within LLMs, rather than cultural or social biases that are studied in the field of LLM safety and fairness [14]., such as learning bias [10], label bias [15], selection biases [16], co-occurrence bias [17], and so forth. In practice, if LLMs establish shortcut ‘FlowersPosition’ based on given samples in Figure 1, predictions may fail when confronted with ‘FlowersNegative’ samples. Apart from this mapping from lexicon to specific labels, shortcuts also exist widely in various forms across different natural language processing tasks. For example, in Natural Language Inference (NLI) [18], Question-Answering (QA) [19], and Machine Reading Comprehension (MRC) [20], LLMs have a preference for overlapping parts of two input branches. These unhealthy shortcuts have been widely proven to be harmful to the robustness [21], generalization [22], safety [23], morality [24], as well as fairness [25], and may lead to hallucinations in LLMs [26].
@@
Given the rapid development of LLMs and the impact of shortcut learning on them, our survey aims to systematically summarize, discuss, and consolidate the existing research, and to sensitize the community to the current advancements. Different from previous reviews on shortcut learning [27, 10, 28, 29], our survey focuses on LLMs, while prior research primarily centered on Lightweight Pre-trained Language Models (LPLMs) such as BERT [30] and RoBERTa [31]. Considering the differences in the training and application of LLMs, we deem it necessary to reassess the types of shortcuts, the causes of shortcut learning, benchmarks, and mitigation strategies, and to highlight the unresolved issues surrounding shortcut learning in LLMs. Additionally, we have summarized different types of shortcuts and their related research tasks, which can assist researchers in clarifying research directions and quickly identifying application scenarios. The organization of sections is as follows.
@@ 2 Types of Shortcuts
Compared to [10], we extend and summarize various types of shortcuts based on the characteristics of NLP tasks in LLMs. Specifically, we classify shortcuts into instinctive shortcuts and acquired shortcuts, based on whether LLMs rely on prompts to establish superficial correlations with labels or answers. But in reality, we believe that instinctive shortcuts can influence acquired shortcuts to some degree. In other words, instinctive shortcuts may facilitate LLMs in acquiring shortcuts from demonstrations [35].
@@ 2.1 Instinctive Shortcuts
Instinctive shortcuts refer to the inherent preferences of LLMs for texts or patterns outside the examples provided in a given context. These preferences are not influenced by the contextual demonstrations but hinder the effective learning of the input-label relationships presented in those demonstrations. To avoid naming ambiguity, we have adopted the label bias definition proposed by [15], which has been widely discussed.
@@
Vanilla-label Bias describes the model’s uncontextual preference for predicting certain label names. Some studies point out that regardless of the prompt, LLMs tend to assign higher probabilities to specific labels [32, 14]. Ari et al. demonstrate the existence of Surface Form Competition among different words, leading to a significant decline in prediction performance when low-frequency words are the correct answers [33]. Similarly, furthur studies also confirm the preference of LLMs for high-frequency labels [34, 17]. Jane et al. show that replacing labels with meaningless words, such as ‘@#$’, leads to a decline in performance compared to using meaningful labels like ‘ABC’ [35] . This preference can sometimes manifest itself in terms of entities, such as LLMs can stubbornly believe that ’Bill Gates’ is the ’founder’ of ’Microsoft’ and ignore the context [36]. This prior preference is harmful in some cases, resulting in the difficulty for LLMs to completely overcome the predicted preference obtained from the pre-training data and consider all contextual information unequally [37].
@@
Context-label Bias reveals the influence of context prompt, including factors such as order and format, on prediction outcomes. For example, subtle adjustments to the prompt format can lead to high instability in model performance, such as changing ‘Passage: text’ to ‘Passage text’ [39]. Moreover, LLMs are also highly sensitive to the order of inputs, whether it be the order of demonstrations or the order of candidate answer options. For example, LLMs are thought to favor options presented at a particular ranking position, such as first or last [40, 41, 14, 16]. Zhao et al. also observe that answers closer to the end of the prompt are more likely to be repeated by the model [32]. Lu et al. demonstrate the sequence of presented samples can lead to a vast difference in performance, ranging from nearly achieving the current state-of-the-art to merely performing at the level of random guessing [38].
@@
Domain-label Bias describes the dependence on prior knowledge related to the task. Research has indicate that the in-domain words randomly sampled from the dataset are heavily biased towards the domain labels in ICL [15]. Accordingly, label words also often imply specific task types [43]. For example, the word ‘positive’ frequently appears in connection with sentiment classification tasks. If ‘positive’ were replaced with a meaningless symbol like ‘1’, LLMs would find it difficult to discern whether the task is sentiment classification or others. This leads to LLMs relying on their semantic priors in ICL predictions, rather than learning from the input-label relationships presented in demonstrations [75]. This prior can also be observed in QA tasks, where LLMs are able to make correct predictions even without being given options [44]. Additionally, LLMs are also highly sensitive to term frequencies and task probabilities [42]. For instance, when asked to decode shift ciphers with various shift levels, GPT-4 scores 50% or above for the three most common shift numbers (1, 3, and 13), but scores below 3% for all other shift levels represented by different numbers [34].
@@
Apart from the semantic priors between texts and labels, entity knowledge within LLMs can also lead to bias [78]. For instance, LLMs may assume that ‘Bill Gates’ is the founder of ‘Microsoft’, even when given the context ‘Bill Gates went to Microsoft Building 99’ where the relationship between ‘Bill Gates’ and ‘Microsoft’ is visitor. Similarly, in relation extraction tasks where the intended relationship to be extracted is unrelated to the existing relationships in the input, LLMs may still extract these unrelated relationships [45].
Figure 4: Examples of different reasoning-label bias. -
@@
Reasoning-label Bias is a newly summarized shortcut type in this paper compared with [15]. It indicates that LLMs tend to get superficial answers directly from the provided examples rather than reasoning the underlying patterns [79] and true causal effect [51], especially in complex reasoning (eg. compositional reasoning [80, 53], logical reasoning [54]) and summarization tasks involving long contexts [46, 47]. Reasoning-label bias is affected by several factors, including the absence and error of inference logic [48], the potential association between the head entity and the entity to be predicted [49, 50], and the order of causal relations in the narrative [52]. For example, in the reasoning example shown in Figure 4, the complex reasoning of multiple hops may cause the reasoning cues of LLMs to be lost. In the case about ‘Olympics’, LLMs may directly build a false inference path that skips ‘Japan’ and still outputs the desired result. At first glance, reasoning-label bias seems to rely on demonstrations, but essentially, it represents a tendency of LLMs to take shortcuts when solving complex problems. Therefore, we categorize it as an instinctive shortcut.
@@ 2.2 Acquired Shortcuts
Lexicon. When spurious correlation are established between certain lexical features and specific labels, the corresponding lexicon may then become a shortcut. Arbitrary text can serve as lexicon shortcuts, including letters, signs, numbers, negation words, adverbs of degree, or even recurring sentences [55]. Among them, negation words have garnered the most attention from researchers, as LLMs often provide contradictory predictions due to these words [57, 58, 59] But studies have indicated that LLMs are more likely to rely on n-grams and content words rather than others, such as stop words [56, 60].
@@
Concept. Concept shortcut refers to the phenomenon where the concepts of certain texts in the demonstrations are disproportionately associated with the labels [12]. For example, if sentences containing cities are associated with negative emotions and samples containing countries are also associated with positive emotions, when LLMs encounter new city words in the test sample, they may predict it as having negative emotions [59].
@@
Overlap. Overlap shortcuts often emerge in tasks that involve two branches, such as NLI and QA. In NLI, overlap between the premise and the hypothesis can be a strong indicator of Entailment [56, 57, 58]. In extractive QA, words that appear in both the question and the background may be used as anchors. LLMs tend to extract answers around these anchors to reduce the error rate [61].
Figure 6: Position Shortcut versus Context-label Bias. -
@@
Position. LLMs use positional information as a basis for prediction. For example, if the answer is located at the beginning or end of the text, the model is less likely to make an error [61, 62]. Conversely, Yu et al. show that adding content unrelated to the answer at the beginning of a sentence can degrade the prediction performance of LLMs [58]. But sometimes, LLMs can briefly read the opening segment to predict the answer without needing to read the entire document [81]. Unlike the positional preference observed in context-label bias, Position refers to the model’s bias resulting from the location of a specific segment of text within the demonstration. The former is caused by the sensitivity of LLMs to the formats of prompts irrespective of the text’s specific content, whereas the latter involves the concrete content of the examples as shown in Figure 6.
@@
Text Style. The writing style of the text is also a possible shortcut [82]. Some studies demonstrate that shifts in presentation style can have an impact on performance, such as when standard English is rewritten into Shakespearean English [55, 58]. LLMs can also capture the correlation between style and labels, leading to a decline in the generalization ability of ICL [59].
@@
Group Dynamics. It refers to the interactions, relationships, and influences among members within a group [83]. Similarly, in the decision-making process, LLMs may evaluate a choice based on other available options. For instance, if other options are even numbers, the model may tend to lean towards an odd number choice [44]. This type of shortcut can also be associated with the A-Not-B error in the field of cognitive science [84]. If ‘A’ appears more frequently in the demonstrations, the model is more inclined to predict ‘A’, even if ‘B’ is correct [32, 63, 64].
@@ 3.1 LLMs Training
Instinctive shortcuts reveal the inherent biases of LLMs. The causes of inherent biases are diverse and closely related to the training process of LLMs, involving factors such as data distribution and instruction tuning. This makes it difficult for LLMs to fully overcome the prediction biases acquired from pre-training data and to consider all contextual information equally during the ICL process [37].
Pretraining. Some studies suggest that instinctive shortcuts originate from the data distribution during the pretraining process [42, 34], which leads to a bias in the predictions of LLMs towards high-frequency and frequently co-occurred words [17, 50]. Furthermore, Yu el al. utilize SHAP analysis [85] to confirm that high-frequency words that have spurious correlations with a certain label contribute more significantly to the model’s predictions for that label [59]. Additionally, if the frequency of a particular LLMs task is low, its prediction results will also be inferior compared to those of high-frequency tasks [34]. Based on the observations mentioned above, further work implicitly assume that there is a shift in the label marginal of the ICL predictive distribution [65, 15], and Jiang et al. theoretically proves that the emergence of shortcuts stems from label shifting caused by the contextual label marginal typically deviating from the true label distribution [67].
In addition to the above effects on the preference for certain labels and words, a biased training corpus can also hamper the inference ability of LLMs. The bias present in the training corpus can cause ICL to rely on irrelevant text ranges in the prompts and generate incoherent chains of thoughts, which impairs the logical reasoning ability of the model [51]. Rare behavioral patterns in the training data, such as converse relations, can also confuse LLMs, causing them to rely on shortcuts when making judgments [66].
Instruction Tuning. Another perspective holds that the instinctive shortcuts in LLMs originate from the process of instruction tuning, during which the LLMs learn spurious correlations within the instructions [57]. In complex multitasking scenarios, LLMs can also learn more intricate associations between tasks, features, and labels during the fine-tuning process. As a result, LLMs can exploit these spurious correlations simply by being provided with useful task information [35]. But research on shortcuts for instruction tuning is still relatively few.